Neglecting air resistance, how long shall it take an object in free fall to have a velocity of 5.0 m/s?




1
Expert's answer
2021-11-15T10:33:20-0500

Solution.

"v_0=0 m\/s;"

"v=5.0 m\/s;"

"g=9.8 m\/s^2;"

"v=v_0+gt\implies t=\dfrac{v-v_0}{g};"

"t=\dfrac{5.0-0}{9.8}=0.5s;"

Answer: "t=0.5 s."